In a titration experiment with your product, it took 1.1 mL of phenylmagnesium bromide to turn the indicator from yellow to orange. What is the concentration of the phenylmagnesium bromide solution if you started with 68.5 mg of indicator? Report you answer to 3 decimal places.
Answer #1

Molar mass of phenylmagnesium bromide = 12*6 + 5 + 24 + 79.9 = 180.9g mol-1

mass of phenylmagnesium bromide taken = 68.5mg

volume taken = 1.1ml

Conc = 68.5*10-3 / 180.9*1.1*10-3 = 0.344M